# SQ-LLaVA: Self-Questioning for Large Vision-Language Assistant

### v1

@@ 3.3 Enhanced Visual Representation
Specifically, we utilize clustering to extract centroid representations of image tokens $Z_{v}$, where each cluster center is treated as a prototype, which, in return, will be distributed to each of the original image token embeddings. Our proposed prototype extractor $\phi(\cdot)$ is a lightweight design involving two parts: 1) cluster center optimization and 2) prototype information distribution. Following Ikotun et al. (2023); Vattani (2009), we randomly initialize $K=256$ cluster centers $C$ and deploy the iterative Expectation-Maximization (EM) clustering process to capture representative semantics in the latent space by
@@ 3.4 Model Training
Unlike text-only LLMs, the general-purpose vision language model also fine-tunes LLM using image tokens as input (see Fig. 2). Therefore, the pre-training stage aims to optimize the LLM by explicitly executing the visual instruction. The proposed SQ-LLaVA adopts Vicuna Zheng et al. (2023) as its instruction LLM, pre-training on massive text corpora to predict the next text token given the previous context, not only containing text but also visual instructions. To achieve this, we organize the pre-training data as $D_{PT}=\{[X^{(1)}_{v},X^{(1)}_{a}],\cdots,[X^{(N)}_{v},X^{(N)}_{a}]\}$, where $N$ is the total number of training samples, and each sample has an image and its related descriptions. Each image and text input pair will be mapped to sequence embeddings ($H_{v}$ and $H_{a}$) as elaborated in Section 3.1. During pre-training, we freeze the vision encoder and LLM and mainly train the prototype extractor $\phi$ and the vision-to-language projector $W$. The pre-training goal is to maximize the probability of the predicted image description $H_{a}$ given an image $H_{v}$. When training a visual instructional LLM, we follow the negative log-likelihood objective function as
@@
Existing methods, such as LLaVA Liu et al. (2023a); Chen et al. (2023), mainly update the vision-to-language projector (usually a couple of linear layers) and the language model during fine-tuning. Nevertheless, the projector might be too weak to capture the relationship between the image and the questions. Following the previous multi-modal understanding method Radford et al. (2021), we unfreeze the vision encoder and LLM during fine-tuning for a joint optimization further to eliminate the gap between the vision and language domain.
To mitigate the heavy computational overhead, we take advantage of LoRA Hu et al. (2022) as a lightweight tuning option that can achieve similar (even better) performance to fully fine-tuning when training large models on a relatively small dataset. We add LoRA in both the vision encoder and LLM. Thus, the learnable parameters $\theta$ of the proposed SQ-LLaVA during fine-tuning represent a combination of all the parameters of LLM-LoRA, ViT-LoRA, prototype extractor $\phi$, and the vision-to-language projector $W$. Given the instruction tuning data $D_{IT}=\{[X^{(1)}_{v},X^{(1)}_{c}],\cdots,[X^{(N)}_{v},X^{(N)}_{c}]\}$, we take the conversational data $X_{c}$ and the image $X_{v}$ as input, mapping them to sequence embedding ($H_{c}$ and $H_{v}$) as elaborated in Section 3.1, and minimize the negative log-likelihood loss for the self-questioning and answering tasks as follows
@@
where $j\in\left\{1,\cdots,P\right\}$, indicating the index of question or answer within the conversational data $X_{c}^{(*)}$. Notably, previous works Liu et al. (2023b); Chen et al. (2023); Zhu et al. (2024) only involve answering tasks, but we introduce visual self-questioning as an additional training task for visual instruction tuning. Eventually, SQ-LLaVA, as a vision-language assistant, not only executes human instructions by optimizing the objective function in Eq. (7) but can raise questions out of the given image after optimizing the Eq. (6), potentially yielding more diverse question-answer guidance and improving multi-modal understanding.
@@ 4.1 Experimental Setting
Our work uses the open-source visual instruction dataset provided by LLaVA Liu et al. (2023b) and ShareGPT4V Chen et al. (2023) for training. Each dataset has large-scale image-text paired data for pre-training and instruction-following data for fine-tuning. Specifically, the instruction data proposed by LLaVA is a comprehensive mixture of COCO, GQA Hudson and Manning (2019), OCR-VQA Mishra et al. (2019), TextVQA Singh et al. (2019), VisualGenome Krishna et al. (2016), RefCOCO Kazemzadeh et al. (2014), and image from InstructBLIP Dai et al. (2023) involves multiple reasoning, spatial understanding, multi-step inference, optical character recognition, and grounding of visual concepts to language. The ShareGPT4V dataset consists of the same mixture image as LLaVA but enrolls more images from other datasets such as SAM Kirillov et al. (2023) and WebData Ordonez et al. (2011); Schuhmann et al. (2021).
[…]
Table 1: Comparison with state-of-the-art methods on ten benchmarks. After training on the same instruction data, SQ-LLaVA traind on Liu et al. (2023b) and $*{}^{*}$ trained on Chen et al. (2023) surpass their baseline model LLaVA-v1.5 and ShareGPT4V on 9 out of 10 and 6 out of 10 benchmarks in the 7B scale, and 8 out of 10 and 6 out of 10 in the 13B scale. The best results are bold and the second-best results are underlined.
[…]
7B InstructBLIP Dai et al. (2023) - 49.2 34.5 60.5 50.1 - 26.2 60.9 36.0 23.7
Qwen-VL Bai et al. (2023) 78.8 59.3 35.2 67.1 63.8 - - - 38.2 7.4
Qwen-VL-chat Bai et al. (2023) 78.2 57.5 38.9 68.2 61.5 - - - 60.6 56.7
LLaVA-v1.5 Liu et al. (2023b) 78.5 62.0 50.0 66.8 58.2 85.9 30.5 63.4 64.3 58.3
ShareGPT4V Chen et al. (2023) 80.6 63.3 57.2 68.4 60.4 86.8 37.6 72.6 68.8 62.2
@@
$*{}^{*}$ 80.3 63.7 55.3 70.5 60.5 87.2 37.6 74.3 66.6 60.0
13B InstructBLIP Dai et al. (2023) - 49.5 33.4 63.1 50.7 78.9 25.6 58.2 - -
LLaVA-v1.5 Liu et al. (2023b) 80.0 63.3 53.6 71.6 58.2 85.9 35.4 70.7 67.7 63.6
ShareGPT4V Chen et al. (2023) 81.0 64.8 55.6 71.2 62.2 - 43.1 79.9 68.5 63.7
@@
$*{}^{*}$ 81.3 65.0 58.2 71.5 61.9 87.4 39.7 80.7 68.5 62.5
Figure 4: Visual self-questioning of SQ-LLaVA-7B. Comparing to the question data provided by GPT4-V (data collected by LLaVA-v1.5 Liu et al. (2023b)), SQ-LLaVA can generate questions with higher diversity i.e. multiple choice and tricky questions.
We also evaluate our model for visual information discovery through captioning. To be specific, we employ testing images from four datasets, i.e., COCO Chen et al. (2015), Flickr30K Young et al. (2014), Conceptual Sharma et al. (2018), and Nocaps Agrawal et al. (2019). For Flickr30K, we adopt the standard split and focus on its test set, which consists of $1000$ samples. We use Nocaps-out-domain split, which exclusively contains images depicting objects absent in COCO. Following previous research Dessì et al. (2023), we use the Nocaps validation set for testing. Following Dessì et al. (2023), we evaluate the proposed method with regular image captioning metrics, e.g., BLEU Papineni et al. (2002) and CIDEr Vedantam et al. (2015). For the open-world methods LLaVA-v1.5 Liu et al. (2023b), ShareGPT4V Chen et al. (2023) and the proposed SQ-LLaVA, we utilize greedy search for caption generation with a prompt of “Provide a brief description of the given image, your answer should be in one sentence.” The generated caption is used to evaluate the performance of image captioning.
4.1.2 Compared Methods.
In this work, we choose baseline models with similar architecture. Precisely, each model consists of a vision encoder and an LLM with different model size and pre-trained weights. For both image description and question answering tasks, we adopt InstructBLIP Dai et al. (2023), Qwen-VL Bai et al. (2023), LLaVA-v1.5 Liu et al. (2023b), and ShareGPT4V Chen et al. (2023) as our baseline models, where all methods were trained on visual instruction data. Besides, we also incorporate ClipCap Mokady (2021) and DiscriTune Dessì et al. (2023), which are trained on regular image-text pair data.
4.1.3 Implementation.
We adopt Vicuna Zheng et al. (2023) as the pre-trained language generative model and CLIP-ViT Radford et al. (2021) as the vision encoder. We pre-train the prototype extractor and the vision-to-language projector using AdamW Loshchilov and Hutter (2019) optimizer with a learning rate of $2\times 10^{-3}$ and a constant scheduler for one epoch. Following previous work Liu et al. (2023b), we keep the global batch size as $256$ for pre-training and $128$ for fine-tuning. During fine-tuning, we insert LoRAHu et al. (2022) with $rank=128$ and $\alpha=256$ into the language model (LLM-LoRA) and LoRA with $rank=32$ and $\alpha=64$ into the vision encoder (ViT-LoRA). We optimize LoRA modules, the prototype extractor, and vision-to-language projector for one epoch. We set the learning rate to $2\times 10^{-4}$ for LoRA, and $2\times 10^{-5}$ for the other layers.
Table 2: Image captioning comparison on four datasets. All methods are trained with images from COCO. SQ-LLaVA is trained on data collected by Liu et al. (2023b). $*{}^{*}$ is trained on data collected by Chen et al. (2023). We adopt B@4 and CIDEr for evaluation.
[…]
B@4 CIDEr B@4 CIDEr B@4 CIDEr B@4 CIDEr
ClipCap Mokady (2021) 2B 8.50 37.03 17.21 41.65 20.32 51.74 1.47 23.74
DiscriTune Dessì et al. (2023) 2B 13.99 53.20 18.48 44.78 24.10 57.06 1.71 28.01
LLaVA-v1.5Liu et al. (2023b) 7B 30.68 112.94 28.67 81.27 35.78 103.56 2.79 39.20
ShareGPT4V Chen et al. (2023) 7B 30.54 110.71 31.00 86.17 37.19 107.45 2.78 37.86
SQ-LLaVA 7B 29.85 110.77 29.88 83.51 36.21 105.42 2.90 39.49
$*{}^{*}$ 7B 30.13 109.71 31.49 83.14 37.20 107.42 2.91 41.24
Figure 5: A qualitative evaluation of detailed and brief image descriptions from three models. We highlight the words and sentences that represent how each model describes the main object in the image.
@@ 4.2 Zero-shot Multilingual Capability
We evaluate SQ-LLaVA on ten benchmarks, covering a range of academic Visual Question Answering (VQA) tasks and recent instruction tuning tasks designed for large vision language models. The academic VQA includes VQA-v2 Goyal et al. (2017) and VizWiz Gurari et al. (2018). GQA Hudson and Manning (2019) is a fine-grained real-world visual reasoning and question-answering benchmark. ScienceQA Lu et al. (2022) is a benchmark with rich subjects (natural science, language science, and social science). TextVQA Singh et al. (2019) requires the model to recognize the texts in the image. LLaVA (in the wild) and MM-Vet Yu et al. (2023) use GPT4 to assess the capability of the models for testing. With manually designed questions, MM-Bench and MMBench-CN Liu et al. (2023c) evaluate the model’s vision-related reasoning and perception for English and Chinese, respectively. POPE Li et al. (2023) is a benchmark for evaluating object hallucination Rohrbach et al. (2018).
@@
In Table 1, we quantitatively compare between SQ-LLaVA and existing models. SQ-LLaVA-7B and SQ-LLaVA-13B trained on two instruction datasets Liu et al. (2023b); Chen et al. (2023) outperform previous methods in six out of ten visual instruction tuning tasks. To be specific, SQ-LLaVA-7B achieves $17.2\%$ improvement over LLaVA-v1.5-7B on the LLaVA (in the wild) benchmark, indicating the superior capabilities of our model in tasks such as detailed description and complex reasoning. Also, SQ-LLavA-7B improves over previous methods on ScienceQA, indicating that our model excels in understanding and reasoning over scientific content and can effectively handle multi-modal information. The improvement in ScienceQA suggests strong capabilities in multi-hop reasoning, comprehension of complex scientific concepts, and the ability to utilize context and explanations to derive correct answers. SQ-LLaVA-7B has a steady improvement over LLaVA-v1.5-7B and ShareGPT4V-7B on the POPE benchmark, and the $2\%$ and $1\%$ improvement indicates that our proposed method has better reliability and trustworthiness since POPE is a task designed to evaluate the phenomenon of object hallucination Li et al. (2023); Rohrbach et al. (2018). In the bottom section of Table 1, the proposed SQ-LLaVA-13B surpasses previous works in six out of ten benchmarks, indicating the scalability of our method on larger LLM. Notably, all the improvements are achieved with significantly fewer trainable parameters compared to other methods Chen et al. (2023); Liu et al. (2023b); Bai et al. (2023). To further evaluate SQ-LLaVA, we present experiments on one recent benchmark in Appendix A.2.
@@ 4.3 Visual Information Discovery
Abilities of SQ-LLaVA Through Qualitative Samples. SQ-LLaVA demonstrates many advanced abilities compared to traditional vision language models. For example, SQ-LLaVA can effectively eliminate object hallucination Li et al. (2023); Rohrbach et al. (2018), delivering the predictions with better trustworthiness. Besides, SQ-LLaVA can generate diverse and meaningful questions about a given image without human textual instructions.
@@
In Fig. 4, we show visual self-questioning results of SQ-LLaVA. We find SQ-LLaVA can sample questions with high diversity and quality. For example, the generated questions include multiple-choice, reasoning, binary choice questions, and the content within each question, all of which are either highly correlated or tricky to the given image, such as “What is the hair color of the man that is looking up at the ocean?” from the third image and “Are there cows in the picture?” from the second image.
@@
The detailed image description example in Fig. 5 demonstrates that SQ-LLaVA can generate visual information with concrete concepts within the image, such as “Hyundai”, “small white dog”, “Nikky Stephen”. In contrast, LLaVA-v1.5 can only describe the image from a general concept, such as “car”, “dog”, or “everyday moment”. Even though ShareGPT4V can generate detailed descriptions, it still suffers from the issue of object hallucination, such as “perhaps waiting” and “behind the duck sign”. By observation, SQ-LLaVA can describe the image with less unintended text, yielding higher reliability. On the other hand, all methods can generate similar brief image descriptions containing only the main objects within the image, indicating the promising summary skills of each model. See Appendix B for more qualitative evaluation on image description and visual self-questioning.
[…]

### v2

@@ 3.3 Enhanced Visual Representation
Specifically, we utilize clustering to extract centroid representations of image tokens $Z_{v}$, where each cluster center is treated as a prototype, which, in return, will be distributed to each of the original image token embeddings. Our proposed prototype extractor $\phi(\cdot)$ is a lightweight design involving two parts: 1) cluster center optimization and 2) prototype information distribution. Following [14, 45], we randomly initialize $K=256$ cluster centers $C$ and deploy the iterative Expectation-Maximization (EM) clustering process to capture representative semantics in the latent space by
@@ 3.4 Model Training
Unlike text-only LLMs, the vision-language model also fine-tunes LLM using image tokens as input (see Fig. 3). Therefore, the pre-training stage aims to optimize the LLM by explicitly executing the visual instruction. The proposed SQ-LLaVA adopts Vicuna [57] as its instruction LLM, pre-training on massive text corpora to predict the next text token given the previous context, not only containing text but also visual instructions. To achieve this, we organize the pre-training data as $D_{PT}=\{[X^{(1)}_{v},X^{(1)}_{a}],\cdots,[X^{(N)}_{v},X^{(N)}_{a}]\}$, where $N$ is the total number of training samples, and each sample has an image and its related descriptions. Each image and text input pair will be mapped to sequence embeddings ($H_{v}$ and $H_{a}$) as elaborated in Section 3.1. During pre-training, we freeze the vision encoder and LLM and mainly train the prototype extractor $\phi$ and the vision-to-language projector $W$. The pre-training goal is to maximize the probability of the predicted image description $H_{a}$ given an image $H_{v}$. When training a visual instructional LLM, we follow the negative log-likelihood objective function as
@@
Existing methods, such as LLaVA [22, 5], mainly update the vision-to-language projector (usually a couple of linear layers) and the language model during fine-tuning. Nevertheless, the projector might be too weak to capture the relationship between the image and the questions. Following the previous multi-modal understanding method [33], we unfreeze the vision encoder and LLM during fine-tuning for a joint optimization further to eliminate the gap between the vision and language domain.
To mitigate the heavy computational overhead, we take advantage of LoRA [12] as a lightweight tuning option that can achieve similar (even better) performance to fully fine-tuning when training large models on a relatively small dataset. We add LoRA in both the vision encoder and LLM. Thus, the learnable parameters $\theta$ of the proposed SQ-LLaVA during fine-tuning represent a combination of all the parameters of LLM-LoRA, ViT-LoRA, prototype extractor $\phi$, and the vision-to-language projector $W$. Given the instruction tuning data $D_{IT}=\{[X^{(1)}_{v},X^{(1)}_{c}],\cdots,[X^{(N)}_{v},X^{(N)}_{c}]\}$, we take the conversational data $X_{c}$ and the image $X_{v}$ as input, mapping them to sequence embedding ($H_{c}$ and $H_{v}$) as elaborated in Section 3.1, and minimize the negative log-likelihood loss for the self-questioning and answering tasks as follows
@@
where $j\in\left\{1,\cdots,P\right\}$, indicating the index of question or answer within the conversational data $X_{c}^{(*)}$. Notably, previous works [21, 5, 58] only involve answering tasks, but we introduce visual self-questioning as an additional training task for visual instruction tuning. Eventually, SQ-LLaVA, as a vision-language assistant, not only executes human instructions by optimizing the objective function in Eq. (7) but can raise questions out of the given image after optimizing the Eq. (6). This capability potentially yields more diverse question-answer guidance and enhances multi-modal understanding.
@@ 4.1 Experimental Setting
Our work uses the open-source visual instruction dataset provided by LLaVA [21] and ShareGPT4V [5] for training. Each dataset has large-scale image-text paired data for pre-training and instruction-following data for fine-tuning. Specifically, the instruction data proposed by LLaVA is a comprehensive mixture of COCO, GQA [13], OCR-VQA [27], TextVQA [39], VisualGenome [17], RefCOCO [15], and image from InstructBLIP [8] involves multiple reasoning, spatial understanding, multi-step inference, optical character recognition, and grounding of visual concepts to language. The ShareGPT4V dataset consists of the same mixture image as LLaVA but enrolls more images from other datasets such as SAM [16] and WebData [29, 36].
We evaluate our model on general vision-language understanding tasks. Specifically, we use ten visual oriented question answering benchmarks: VQAv2 [10]; GQA [13]; VizWiz [11]; SQAI [25]: ScienceQA-IMG; VQAT: TextVQA [39]; POPE [20]; MM-Vet [53]; LLaVAW [22]: LLaVA (in the wild); MMB: MMBenchmark [23]; MMBCN: MMBench-Chinese [23]. We report the prediction accuracy on all the benchmarks. We also evaluate our model for visual information discovery through captioning. To be specific, we employ testing images from four datasets, i.e., COCO [6], Flickr30K [52], Conceptual [37], and Nocaps [2]. Following [9], we evaluate the proposed method with regular image captioning metrics, e.g., BLEU [32] and CIDEr [46]. For the open-world methods LLaVA-v1.5 [21], ShareGPT4V [5] and the proposed SQ-LLaVA, we utilize greedy search for caption generation with a prompt of “Provide a brief description of the given image, your answer should be in one sentence.” The generated caption is used to evaluate the performance of image captioning.
4.2 Implementation
We adopt Vicuna [57] as the pre-trained language generative model and CLIP-ViT [33] as the vision encoder. We pre-train the prototype extractor and the vision-to-language projector using AdamW [24] optimizer with a learning rate of $2\times 10^{-3}$ and a constant scheduler for one epoch. Following previous work [21], we keep the global batch size as $256$ for pre-training and $128$ for fine-tuning. During fine-tuning, we insert LoRA[12] with $rank=128$ and $\alpha=256$ into the language model (LLM-LoRA) and LoRA with $rank=32$ and $\alpha=64$ into the vision encoder (ViT-LoRA). We optimize LoRA modules, the prototype extractor, and vision-to-language projector for one epoch. We set the learning rate to $2\times 10^{-4}$ for LoRA, and $2\times 10^{-5}$ for the other layers. All the weights of the pre-trained language model and vision encoder remain fixed during fine-tuning.
Table 1: Comparison with state-of-the-art methods on ten benchmarks. After training on the same instruction data, SQ-LLaVA traind on [21] and SQ-LLaVA trained on [5] surpass their baseline model LLaVA-v1.5 and ShareGPT4V on 9 out of 10 and 6 out of 10 benchmarks in the 7B scale, and 8 out of 10 and 6 out of 10 in the 13B scale. The best results are bold and the second-best results are underlined.
LLM Model VQAv2 GQA VizWiz SQAI VQAT POPE MM-Vet LLaVAW MMB MMBCN
7B InstructBLIP [8] - 49.2 34.5 60.5 50.1 - 26.2 60.9 36.0 23.7
Qwen-VL [3] 78.8 59.3 35.2 67.1 63.8 - - - 38.2 7.4
Qwen-VL-chat [3] 78.2 57.5 38.9 68.2 61.5 - - - 60.6 56.7
LLaVA-v1.5 [21] 78.5 62.0 50.0 66.8 58.2 85.9 30.5 63.4 64.3 58.3
ShareGPT4V [5] 80.6 63.3 57.2 68.4 60.4 86.8 37.6 72.6 68.8 62.2
@@ 4.2 Implementation
SQ-LLaVA 80.3 63.7 55.3 70.5 60.5 87.2 37.6 74.3 66.6 60.0
13B InstructBLIP [8] - 49.5 33.4 63.1 50.7 78.9 25.6 58.2 - -
LLaVA-v1.5 [21] 80.0 63.3 53.6 71.6 58.2 85.9 35.4 70.7 67.7 63.6
ShareGPT4V [5] 81.0 64.8 55.6 71.2 62.2 - 43.1 79.9 68.5 63.7
@@
SQ-LLaVA 81.3 65.0 58.2 71.5 61.9 87.4 39.7 80.7 68.5 62.5
Figure 5: Visual self-questioning of SQ-LLaVA-7B. Comparing to the question data provided by GPT4-V (data collected by LLaVA-v1.5 [21]), SQ-LLaVA can generate questions with higher diversity i.e. multiple choice and tricky questions.
@@ 4.3 Zero-shot Multilingual Capability
We evaluate SQ-LLaVA on ten benchmarks, covering a range of academic Visual Question Answering (VQA) tasks and recent instruction tuning tasks designed for large vision language models. The academic VQA includes VQA-v2 [10] and VizWiz [11]. GQA [13] is a fine-grained real-world visual reasoning and question-answering benchmark. ScienceQA [25] is a benchmark with rich subjects (natural science, language science, and social science). TextVQA [39] requires the model to recognize the texts in the image. LLaVA (in the wild) and MM-Vet [53] use GPT4 to assess the capability of the models for testing. With manually designed questions, MM-Bench and MMBench-CN [23] evaluate the model’s vision-related reasoning and perception for English and Chinese, respectively. POPE [20] is a benchmark for evaluating object hallucination [34].
@@
In Table 1, we quantitatively compare between SQ-LLaVA and existing models. SQ-LLaVA-7B and SQ-LLaVA-13B trained on two instruction datasets [21, 5] outperform previous methods in six out of ten visual instruction tuning tasks. To be specific, SQ-LLaVA-7B achieves $17.2\%$ improvement over LLaVA-v1.5-7B on the LLaVA (in the wild) benchmark, indicating the superior capabilities of our model in tasks such as detailed description and complex reasoning. Also, SQ-LLavA-7B improves over previous methods on ScienceQA, indicating that our model excels in understanding and reasoning over scientific content and can effectively handle multi-modal information. The improvement in ScienceQA suggests strong capabilities in multi-hop reasoning, comprehension of complex scientific concepts, and the ability to utilize context and explanations to derive correct answers. SQ-LLaVA-7B has a steady improvement over LLaVA-v1.5-7B and ShareGPT4V-7B on the POPE benchmark, and the $2\%$ and $1\%$ improvement indicates that our proposed method has better reliability and trustworthiness since POPE is a task designed to evaluate the phenomenon of object hallucination [20, 34]. In the bottom section of Table 1, the proposed SQ-LLaVA-13B surpasses previous works in six out of ten benchmarks, indicating the scalability of our method on larger LLM. Notably, the performance inconsistency on some datasets might be due to the unsupervised prototype extractor (lacking pixel-wise guidance) in our model. To mitigate this issue, we could leverage pseudo object masks (e.g., given by the pre-trained segment anything) in learning prototypes. Despite this limitation, all the improvements are achieved with significantly fewer trainable parameters compared to other methods [5, 21, 3].
Table 2: Comparison of image captioning on four datasets. SQ-LLaVA is trained on data collected by [21]. SQ-LLaVA is trained on data collected by [5].
Model Flickr30k Nocapsout Conceptual
B@4 CIDEr B@4 CIDEr B@4 CIDEr
ClipCap [28] 17.21 41.65 20.32 51.74 1.47 23.74
DiscriTune [9] 18.48 44.78 24.10 57.06 1.71 28.01
LLaVA-v1.5[21] 28.67 81.27 35.78 103.56 2.79 39.20
ShareGPT4V [5] 31.00 86.17 37.19 107.45 2.78 37.86
SQ-LLaVA 29.88 83.51 36.21 105.42 2.90 39.49
SQ-LLaVA 31.49 83.14 37.20 107.42 2.91 41.24
(a) Zero-shot image captioning
Model Zero-shot Fine-tune
B@4 CIDEr B@4 CIDEr
ClipCap 8.50 37.03 32.60 108.55
DiscriTune 13.99 53.20 32.31 105.40
eP-ALM [38] 29.47 97.22 33.35 111.63
MAPL [26] 12.30 54.30 36.45 125.20
LLaVA-v1.5 29.96 111.46 - -
SQ-LLaVA 29.85 110.77 40.76 136.78
(b) Image captioning on COCO
@@ 4.4 Visual Information Discovery
Abilities of SQ-LLaVA Through Qualitative Samples. SQ-LLaVA exhibits numerous advanced capabilities compared to traditional vision-language models. Notably, SQ-LLaVA effectively mitigates object hallucination [20, 34], resulting in predictions that are more trustworthy. Additionally, SQ-LLaVA is capable of generating diverse and meaningful questions about a given image without requiring human textual instructions. As illustrated in Fig. 5, the generated questions include multiple-choice, reasoning, binary choice questions, and the content within each question, all of which are either highly correlated or tricky to the given image, such as “What is the hair color of the man that is looking up at the ocean?” from the third image and “Are there cows in the picture?” from the second image. These examples highlight SQ-LLaVA’s ability to produce questions with high diversity and quality.
@@
The detailed image description example in Fig. 6 demonstrates that SQ-LLaVA can generate visual information with concrete concepts within the image, such as “Hyundai”, “small white dog”, “Nikky Stephen”. In contrast, LLaVA-v1.5 can only describe the image from a general concept, such as “car”, “dog”, or “everyday moment”. Even though ShareGPT4V can generate detailed descriptions, it still suffers from the issue of object hallucination, such as “perhaps waiting” and “behind the duck sign”. By observation, SQ-LLaVA can describe the image with less unintended text, yielding higher reliability.
Figure 6: A qualitative evaluation of detailed image descriptions from three models. We highlight the words and sentences that represent how each model describes the main object in the image.
Quantitative Analysis. SQ-LLaVA serves as a general-purpose vision-language model and enables zero-shot image captioning. As indicated in Table 2(a), SQ-LLaVA achieves $73\%$ and $66\%$ averaged improvement over ClipCap and DiscriTune on all datasets, indicating the effectiveness of visual instruction tuning. Compared with the baseline model LLaVA-V1.5 [21], SQ-LLaVA achieves $2\%$ averaged improvement on all datasets, indicating the effectiveness of visual self-questioning. Also, we find SQ-LLaVA surpasses the baseline model ShareGPT4V [5] on Nocapsout and Conceptual dataset, demonstrating the adaptability of the proposed method on unseen testing images from new domains. Moreover, as shown in Table 2(b), SQ-LLaVA can be easily adapted to COCO captioning via instruction tuning on short descriptions.
4.5 Ablation Study
In Table 3, we conduct experiments with different architecture designs and training strategies on five question-answering benchmarks. For a fair comparison, we train the baseline models on our local machine with the same training recipe of LLaVA-LoRA [21]. Specifically, we present our full model and three ablated models by removing one component each time. We adopt the dataset [21] with $558$k for pre-training (PT) and $665$k for fine-tuning (FT). As compared, self-questioning (SQ) brings a consistent performance boost on all the five benchmarks, indicating the effectiveness of visual self-questioning on improving visual language understanding. Besides, we introduce the prototype extractor (Proto) to enhance visual representation, achieving $0.9\%$ improvement in average accuracy among five benchmark. With all three components incorporated, we observe a $2.4\%$ improvement in average accuracy.
As shown by the bottom block of Table 3, we conduct experiments with the same ablation settings but with a larger scale of the visual instruction data [5] (i.e., for both PT and IT). Overall, SQ-LLaVA achieves $2.4\%$ improvement over the baseline model after training on the smaller dataset and achieves $3.0\%$ improvement after training on the larger dataset.